# Robustly optimal dynamics for active matter reservoir computing

### v1

@@ IV Discussion and Conclusion
We conducted an empirical simulation study of reservoir computing using a physical model system composed of active matter agents, driven chaotically via local repulsive forces. In comparison to earlier work [63], we discovered an exceptional dynamical regime of the reservoir that appears to be robustly optimal for computation across various circumstances.
We find that the parameters that regulate the microscopic particle speed directly, via non-Hamiltonian forces at a microscopic scale, are the most influential physical controls: They stand as lowest-order dynamical causes for attaining high computational performance; they further enable the system to adapt and render high performances in various controlled contexts. We infer that collective effects (via particle repulsions) must be of the next order, as a comparison between the one- and two-particle reservoirs versus the full $N=200$ systems shows a boost of around 30 % performance. The effects of non-Hamiltonian particle-particle alignment forces (as defined in this model) appear to be of higher order; in the optimal computing regime, these can be eliminated without harming performance.
Therefore, the system’s adaptability to significant environmental fluctuations, which are induced by the driver, originates at a microscopic level and should hence be generically detectable. We investigated the intrinsic relaxation characteristics of noninteracting (or, likewise, interacting) agents that are free from external driving forces: In an ensemble starting from an initial state, the mean microscopic relaxation dynamics exhibit a rapid exponential decay, followed by a brief power-law decay, eventually transitioning to a final steady state characterized by weak oscillations.
This microscopic dynamical origin of adaptivity seems to guarantee that optimal computing is largely independent of the specific driving pattern or the inference task at hand; we tested this empirically. There is one caveat: the driver speed should somehow stay within the same order-of-magnitude speed of the agents, or on that of propagating acoustic signals. Intriguingly, one can observe how this kind of adaptability manifests in an efficient erasure of long-term information while gradually transferring and transforming relevant dynamical information during a proper time window – seen phenomenologically through the simulation videos, but also quantitatively in the velocity autocorrelation function.
Also previously alluded to by a comment in Ref. [63], canonical (mean) order parameters (polarity, rotation) indicating the global order of a swarm, do not capture the essence of adapting to external changes, thus, do not indicate areas of high predictive performance in general. In contrast, our approach demonstrated that correlation functions on agent dynamics (i.e. correlated velocity fluctuations and velocity autocorrelations in time) do so, and offer useful information for interpreting the dynamics generating adaptive behavior.
@@
Our results suggest the emergence of interfaces may be fundamental to the functionality of computing systems in soft matter (of analog or neuromorphic kinds): The interface that forms around the driver appears to be a crucial feature for the computational properties of the RC. It might seem counterintuitive, as a slowly moving interface exhibits little variation, meaning there are few significant fluctuations informative for the prediction task. Indeed, Ref. [63] comments on “adiabatic” driving and the resulting difficulty for making a prediction. However, a coherent state enables sharp detections: Specifically, transitions and fluctuations to and from the highly nonlinearly responding state reported in Sect. III.3 will be readable from signals in the observation layer (a follow-up study investigating this is underway). From another view, surface phenomena are known to be rich [110], particularly under nonequilibrium conditions [111, 112, 113], which may suggest a kind of capacity for “physical expressivity” – to amplify or express subtle changes in the physical system at lower scales [80]. For the Lorenz-63 driving, near the quasi-stationary state, the radial propagation of velocity fluctuations originating at the interface is clearly visible, as is the propagation of fluctuations during shear thinning when the driver moves through the medium during the highly nonlinear response. The integration of the driver within the soft matter system facilitates a continuous transfer of information from any changes in the driver’s state into the system.
@@
We speculate that in the cases of simpler tasks, more of the information processed (as defined via the Gaussian kernels that measure local densities in structure and momentum) expresses structural changes, which are inherently slower variables and thus have more internal coherence. In other words, the regular patterns over space and time-scales are detectable in a larger amount in structural fluctuations. This is subject to future investigation. For these simpler tasks, the dynamic quantifiers of the responding system seem to be less clear indicators of information processing abilities (in the more regularly changing environments) – at least in the way they were defined in this study. Quantifiers taking into consideration different dynamical symmetries (i.e., macroscopic rotations) would clarify this outstanding issue.
@@
Representation: To take the above statements a step further, we reckon that the various modes or patterns of dynamical and structural fluctuations are essential to the informational capacity of the system, i.e., the range of information available to the final RC layer, which is then transformed with a linear model. Further, the complexity of the emergent mechanisms in the system is informative about the nature of the driving forces and how difficult the resulting task is for the system to represent. For example, whether some mechanisms appear nested within others, are decomposable [114], or entail redundancies [115], are signs (which also pertain to robustness and the degree of nonlinearity of couplings manifesting in the nonequilibrium system, which we discuss below).
Robustness, approximation: Notably, it is thought that this adaptivity often comes with a tradeoff in robustness [116]; we claim somewhat the opposite: a system that is adaptive is also robust – in a “biological” analogy that it can continue to render useful functions despite major intrinsic changes or constraints [117, 115]. One conceptual subtlety may be that locally in space and time, the specific response mechanisms and resulting spatiotemporal patterns could change drastically depending on the signals from the current environment (i.e. the driver’s local fluctuations), thus making the system appear “non-robust”. The robustness of this kind might be linked to the approximation property of RC [28]: sufficiently similar signals will lead to similar outputs, as only drastic changes in the environment are amplified.
@@
Within the machine learning literature, the concept of robustness is defined in various ways [118]. As discussed in the reference, it may be crucial to test a system across different models, rather than focusing solely on datasets. In contrast, how robustness is expressed in physical systems is still in the early stages of research [55, 119]. (It has been theorized that it is deeply related to metastability [81]). An understanding could involve the response to internal structural and dynamical changes, confinement, the occlusion of particular higher-level interactions that emerge in the system, and the introduction of random noise to the signals within the system. All of these factors were considered in our tests.
Nonlinearity: How much nonlinearity is needed, and of what kind [46]? An answer may be nuanced, as a complex system renders both reversible and irreversible processes, and, additionally, depending on what features one observes, this translates into different degrees of nonlinearity. The phenomenology (Sect. III.3) reveals a role of quasi-stationarity in parts of the response, and exactly how much of linear as well as nonlinear terms are inherent to the information processing is currently unclear [77, 120].
Time-delay: A basic idea about the origins of computational abilities is time-delay response [121, 122, 41, 40]. This will be related to the nonlinearity in dynamics, but the question is at which scale of treatment the effects originate. We showed that delay can be seen visually at the level of single-particle reservoirs, and is first revealable in the intrinsic relaxation dynamics.
Separation property: Separating different input trajectories into distinct patterns is a key feature of functioning reservoirs [123, 28]. From our findings, separation may be about effectively mapping the driver dynamic fluctuations into those of the responding systems. This may have to involve the proper amplification of contrast that originates at a microscopic scale in dynamics (locally in space and time).
Fading memory: For the optimal dynamic regime, the swarm system shows pronounced signatures of exponential decay at short delay times in its VAC. Viscous motion is inherent, and among other effects, plays a prominent role during the self-healing process, returning the system to the quasi-stationary state (particularly for the Lorenz-63 system). Viscous motion is obviously involved in the erasure of recent “memory” of the driver’s strong recent fluctuations. By its very nature, it is irreversible. The nature of physical memory [124] can be illuminated, e.g., by intrinsic relaxation dynamics. Clearly, the more overdamped the dynamics, the more irreversibility will be inherent to the reservoir’s internal functioning. Yet, the intermediate power-law-like regime seen in the dynamical relaxation near critical damping appears associated with irreversibility arising from self-propelled directed motion.
Collectivity, analog computing: As we illustrated, the factors contributing to optimal performance appear to be general and independent of specific computational tasks; it would be helpful to explain this in more generic terms of collectivity and computation. Outside of computing, it has been observed that collective responsivness leads to intelligent and adaptive behavior [125, 114, 126, 127]. As a first, generalization abilities of information processing systems arise from a final aggregation of information [114], whereby this information needs two specific features: amplification combined with decomposability [125]. The strong diversity of phenomenological mechanisms that arose, particularly visible in the Lorenz-63 system–in the form of quasi-stationary and nonlinear responding states – attests to one form of decomposition. Moreover, the initial exponential decay in the VAC, followed by harmonic-like peaks, exposes partly disentangled dynamical time-scales and mechanisms. The strength of the peak and depth of the minimum of the CVC, as well as the dynamic susceptibility, all attest to the amplification of information occurring in near-critically damped systems. Secondly, there is an intriguing bridge to computation: nesting irreversible components of information processing into reversible ones is part of the design of analog computers [71]. The collective mechanisms during computation in this active matter system exhibited elastic-like (thus coherent, reversible) responsiveness combined with highly viscous (thus incoherent, irreversible) behavior. In light of the RC fading memory property and our results, we reckon that self-restoration may be another feature, and will include “forgetting” [124]; memory erasure is an inevitable thermodynamic property of computation [69].
@@
Computing near “criticality”: A broadly cited hypothesis is that systems poised near criticality generally exhibit adaptive behavior when faced with changing environments [50, 128, 48]. In such contexts, “criticality” often refers to a system-wide nonequilibrium phase transition. In our work, the concept of “critical damping” refers to the absence of oscillations in microscopic dynamics, which alters the overall dynamical behavior of the system. Though one can analyze the scaling behavior in greater detail, a further differentiation of the term “edge of criticality” may be helpful [52].
@@
Interpretability: To make a point on the inherent interpretability offered by information processing machines made out of matter: To contrast, standard employed artificial neural networks present significant challenges to understanding the network’s overall functional operation and internal mechanisms [54, 57]. Understanding how to compute with physical systems requires a paradigm shift, formulating it in terms of dynamical processes rather than algorithms [13, 1] – physics inherently offers explanatory frameworks for dynamics.
@@
From our above discussion on the order-of-importance of types of forces in the model, we envision seeking the truly global optimum for performance. Our next step will be to fine-tune the interparticle repulsion forces (which need not necessarily be long-ranged) and the driver-agent interaction (which may be task dependent). We aim to find further deep connections between the physical properties of the reservoir substrate and its computational properties. We have preliminarily investigated much larger systems ($N=500\dots 10^{5}$) and foresee the different scaling of agent number (and observation kernel) densities to predictive performance in different dynamical regimes. In a further step, other types of higher-order velocity-dependent interparticle forces could potentially enhance performance. Finally, we are studying the effects of additional “artificial” or extrinsically-controllable interactions, such as those for the phase coupling of mobile oscillators (swarmalators), of interest for robotic or AI-agent systems [129, 130, 131]. Their order of importance is unclear at this point and will depend strongly on the specific implementation.
Capturing the right, relevant collective variables in the observation layer poses another outstanding question. The kernels’ distribution in space and scale is one thing to consider [63], the way they compute an average quantity internally is another. The physically relevant information they extract remains conceptually opaque at a deeper level [57].
@@ Acknowledgements.
Funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2075 – 390740016. We acknowledge the support of the Stuttgart Center for Simulation Science (SimTech), the Interchange Forum for Reflecting on Intelligent Systems (IRIS) at the University of Stuttgart, the International Max Planck Research School for Intelligent Systems (IMPRS-IS), and the WIN-Kolleg of the Heidelberg Academy of Science and the Humanities. We thank Christian Holm, Mathias Niepert, Anna Levina, Francesco Ginelli, Max Weinmann, Patrick Egenlauf, Gaurav Gardi, and Lasse J. Schulz for fruitful discussions.
@@ Appendix A Chaotic Time-Series Benchmarking Details
In this paper, we use some common chaotic attractors as benchmark systems for the time series prediction task of our active matter reservoir computer (see Sect. II.2). In our default implementation, we integrate the coupled differential equations of the Lorenz-63 system [86] ($\dot{x}=a(y-x);\dot{y}=x(b-z)-y;\dot{z}=xy-cz$) with standard parameters ($a=10;b=28;c=8/3$) using Euler integration with an integration time step $\Delta t=0.02$ and initial conditions $(x_{0},y_{0},z_{0})=(0.0,1.0,1.05)$ for training and $(0.0,1.1,-1.5)$ for testing. For the other chaotic attractors (Hénon-Heiles, Rössler, Chua, Lorenz-96) we generate trajectories using the dysts library (version 0.6) [84] (corresponding publications: Refs. [83, 31]). To compare predictive performance, we choose to normalize these time series to the mean effective driver speed
@@ Appendix C Data Availability
All data in this manuscript was generated using the ResoBee software for active matter reservoir computing. It is currently being prepared for publication in an open-source software journal.
@@ D.4 Intrinsic Agent Relaxation Dynamics
Figure 20: Intrinsic agent relaxation dynamics for simulations without (a,b) and with (c-f) agent-agent interactions in place and with different integration time steps analogously to Fig. 3a,b. The smaller integration time step of $\Delta t=0.002$ allows resolving relaxation dynamics in the far overdamped regime (nabla symbol) that is otherwise rendered as an arrested regime for $\Delta t=0.02$. The presence of agent-agent interactions (d,f) (which also applies to (c,e)) prominently hinders the oscillations in the relaxation dynamics above critical damping (blue square, purple diamond, black cross), and subtly shifts the time-points of relaxation for the near-critical and overdamped regimes (blue-green triangle, green circle). Visible in (f), the minimally underdamped cases (blue square, purple diamond) morph towards critical or over-damping cases via the collisional and alignment interparticle interactions.

### v2

@@ IV Discussion and Conclusion
We conducted an empirical simulation study of reservoir computing using a physical model system composed of active matter agents, driven chaotically via local repulsive forces. In comparison to earlier work [62], we discovered an exceptional dynamical regime of the reservoir that appears to be robustly optimal for computation.
We find that the parameters that regulate the microscopic particle speed directly, via non-Hamiltonian forces at a microscopic scale, are the most influential physical controls: They stand as lowest-order dynamical causes for attaining high computational performance; they further enable the system to adapt and render high performances in various controlled contexts. We infer that collective effects (via particle repulsions) must be of the next order, as a comparison between the one- and two-particle reservoirs versus the full $N=200$ systems shows a boost of around 30 % performance. Details of driver-agent interactions play a key role as well. The effects of velocity-dependent particle-particle alignment forces (as defined in this model) appear to be of higher order; in the optimal computing regime, these can be eliminated without harming performance.
The substrate’s responsiveness and adaptability can be understood in the sense that it both detects changes in its environment (the driver’s fluctuations) and efficiently self-restores itself from recent history. As these abilities appear to originate at the microscopic level, they are generic. Investigated were the intrinsic relaxation characteristics of noninteracting (or, likewise, interacting) agents that are free from external driving forces: In an ensemble starting from an initial state, the mean microscopic relaxation dynamics exhibit a rapid exponential decay, followed by a brief power-law decay, eventually transitioning to a final steady state characterized by weak oscillations.
This microscopic dynamical origin of these abilities guarantees that optimal computing is largely independent of the specific driving pattern or the inference task at hand. We tested this empirically through four different chaotic driving protocols. We note one caveat: the driver speed should probably stay within the same order-of-magnitude speed of the agents, but this would have to be examined in more detail elsewhere.
As previously noted in Ref. [62], canonical (mean) order parameters (polarity, rotation) indicating the global order of a swarm, do not capture the essence of adapting to external changes and thus cannot reveal areas of high predictive performance in general. Our approach demonstrated that correlation functions on agent dynamics (i.e. correlated velocity fluctuations and velocity autocorrelations in time) capture the responsiveness of an active matter reservoir substrate to its external perturbation and could serve as a more reliable indicator.
@@
Our results suggest the emergence of interfaces may be fundamental to the functionality of computing systems in soft matter (of analog or neuromorphic kinds): The interface that forms around the driver appears to be a crucial feature for the computational properties of the RC. It might seem counterintuitive, as a slowly moving interface exhibits little variation, meaning there are few significant fluctuations informative for the prediction task. Indeed, Ref. [62] comments on “adiabatic” driving and the resulting difficulty for making a prediction. However, a coherent state enables sharp detections: Specifically, transitions and fluctuations to and from the highly nonlinearly responding state reported in Sect. III.3 can be read from signals in the observation layer. From another view, surface phenomena are known to be rich [112], particularly under nonequilibrium conditions [113, 114, 115], which may suggest a kind of capacity for “physical expressivity” – to amplify or express subtle changes in the physical system at lower scales [82]. For the Lorenz-63 driving, near the quasi-stationary state, the radial propagation of velocity fluctuations originating at the interface is clearly visible, as is the propagation of fluctuations during shear thinning when the driver moves through the medium during the highly nonlinear response. The integration of the driver within the soft matter system facilitates a continuous transfer of information from any changes in the driver’s state into the system.
@@
We speculate that in the cases of simpler tasks, more of the information processed (as defined via the Gaussian kernels that measure local densities in structure and momentum) expresses structural changes, which are inherently slower variables and thus have more internal coherence. This is subject to future investigation. For these simpler tasks, the dynamic quantifiers of the responding system seem to be less clear indicators of information processing abilities (in the more regularly changing environments) – at least in the way they were defined in this study. Quantifiers taking into consideration different dynamical symmetries (i.e., macroscopic rotations) would clarify this outstanding issue.
@@
Robustness: Notably, it is thought that sensitivity can come with a tradeoff in robustness [116]. Yet, a system that is adaptive is also robust: For example, a biological cell can continue to render useful behavior despite major intrinsic or external changes [117, 118]. Within the machine learning literature, the concept of robustness is defined in various ways [119]. As discussed in the reference, it may be crucial to test a system across different models, rather than focusing solely on datasets. In contrast, how robustness is expressed in physical systems is still in the early stages of research [54, 120]. It may be associated with metastability [83] or other nonequilibrium features expressing self-restoration. Tests we considered in our study varied the intrinsic physical parameters and inference tasks, which represent changes in static confinement, a new dynamical environment, the occlusion of particular internal interactions, and the introduction of random noise to the system.
Collectivity, analog computing: Outside of computing, it has been observed that collectivity can lead to intelligent and adaptive behavior [121, 122, 123, 124]. The collective mechanisms during computation in this active matter system exhibited a complex mixture of elastic-like and viscous properties. There may be an intriguing bridge to computation: Nesting irreversible components of information processing into reversible ones is part of the design of analog computers [71].
@@
Computing near “criticality”: A broadly cited hypothesis is that systems poised near criticality generally exhibit adaptive behavior when faced with changing environments [125, 126, 127, 128, 129]. In such contexts, “criticality” often refers to a system-wide nonequilibrium phase transition. In our work, the concept of “critical damping” refers to the absence of oscillations in microscopic dynamics, which alters the overall dynamical behavior of the system. Though one can analyze the scaling behavior in greater detail, a further differentiation of the term “edge of criticality” may be helpful [130].
@@
Interpretability: The inherent interpretability offered by information processing machines made out of matter is interesting for stimulating scientific inquiry. Standard employed artificial neural networks present significant challenges to understanding the network’s overall functional operation and internal mechanisms [53, 56]. For physical computing systems, dynamical processes take precedence over algorithmic descriptions [13, 1], and physics provides intrinsic frameworks for understanding such dynamics.
@@
From our above discussion on the order-of-importance of types of forces in the model, we envision seeking the truly global optimum for performance. Our next step will be to fine-tune the interparticle repulsion forces (which need not necessarily be long-ranged) and the driver-agent interaction. We aim to find further deep connections between the physical properties of the reservoir substrate and its computational properties. We have preliminarily investigated much larger systems ($N=500\dots 10^{5}$) and foresee the different scaling of agent number (and observation kernel) densities to predictive performance in different dynamical regimes. In a further step, other types of higher-order velocity-dependent interparticle forces could potentially enhance performance. Finally, we are studying the effects of additional “artificial” or extrinsically-controllable interactions, such as those for the phase coupling of mobile oscillators (swarmalators), of interest for robotic or AI-agent systems [131, 132, 133].
Capturing the right, relevant collective variables in the observation layer poses another outstanding question. The kernels’ distribution in space and scale is one thing to consider [62], the way they compute an average quantity internally is another.
@@ Acknowledgements.
Funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2075 – 390740016. We acknowledge the support of the Stuttgart Center for Simulation Science (SimTech), the Interchange Forum for Reflecting on Intelligent Systems (IRIS) at the University of Stuttgart, the International Max Planck Research School for Intelligent Systems (IMPRS-IS), and the WIN-Kolleg of the Heidelberg Academy of Science and the Humanities. We thank Christian Holm, Mathias Niepert, Anna Levina, Francesco Ginelli, Shankar P. Das, Max Weinmann, Patrick Egenlauf, Gaurav Gardi, and Lasse J. Schulz for fruitful discussions.
@@ Appendix A Chaotic Time-Series Benchmarking Details
In this paper, we use some common chaotic attractors as benchmark systems for the time series prediction task of our active matter reservoir computer (see Sect. II.2). In our default implementation, we integrate the coupled differential equations of the Lorenz-63 system [88] ($\dot{x}=a(y-x);\dot{y}=x(b-z)-y;\dot{z}=xy-cz$) with standard parameters ($a=10;b=28;c=8/3$) using Euler integration with an integration time step $\Delta t=0.02$ and initial conditions $(x_{0},y_{0},z_{0})=(0.0,1.0,1.05)$ for training and $(0.0,1.1,-1.5)$ for testing. For the other chaotic attractors (Hénon-Heiles, Rössler, Chua, Lorenz-96), we generate trajectories using the dysts library (version 0.6) [86] (corresponding publications: Refs. [85, 31]). To compare predictive performance, we choose to normalize these time series to the mean effective driver speed
@@ Appendix C Data Availability
All data in this manuscript were generated using the ResoBee software for active matter reservoir computing. It is currently being prepared for publication in an open-source software journal.
@@ D.4 Intrinsic Agent Relaxation Dynamics
Figure 20: Intrinsic agent relaxation dynamics for simulations without (a,b) and with (c-f) agent-agent interactions in place and with different integration time steps analogously to Fig. 3a,b. The smaller integration time step of $\Delta t=0.002$ allows resolving relaxation dynamics in the far overdamped regime (nabla symbol) that is otherwise rendered as an arrested regime for $\Delta t=0.02$. The presence of agent-agent interactions (d,f) (which also applies to (c,e)) prominently hinders the oscillations in the relaxation dynamics above critical damping (square, diamond, cross), and subtly shifts the time-points of relaxation for the near-critical and overdamped regimes (triangle, circle).